A flat surface of area 1.25 m2 is rotated through a uniform horizontal electric field of 5 N/C. What is the electric flux if the surface is (a) parallel and (b) perpendicular to the electric field?





1
Expert's answer
2022-02-23T12:11:10-0500

Answer

Area

A="1.25m^2"

Electric field E=5N/C


Electric flux is given by

"\phi=EA\cos\theta"

a) when parallel then angle is 0°

So cos °=1

So flux

"\phi=5*1.25*1=6.25Wb"

b) when perpendicular then

Angle is 90° so

Cos 90°=0

Then flux will be zero.